OpenStudy (anonymous):

p^3-4p+9 is a perfect square number find the value of p, for p is positive integer

OpenStudy (anonymous):

@dan815

OpenStudy (jdoe0001):

heheh, for a perfect square, looks very "cubic" \(\bf p^{\color{brown}{ 3?}}-4p+9\)

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

yeah, that's p cubic not p square

geerky42 (geerky42):

Where is \(n\) in \(p^3-4p+9\)?

OpenStudy (jdoe0001):

yeah... the "n" came out of the woodwork, noticed that one

OpenStudy (anonymous):

oppss.. sorry that should to find value of p p^3 -4p + 9 = k^2 i was trying to find value of p, but some method is loss :)

geerky42 (geerky42):

Funny, this is kind of beyond my knowledge, but by brutal force, I just discovered that p=2 works.

geerky42 (geerky42):

Because \(p^3-4p = (p+2)p(p-2)\), so let p=2 and we have \(p-2=0\) so we are left with \(9\), which is square number.

geerky42 (geerky42):

maybe this isn't what you are looking for?

OpenStudy (jdoe0001):

hmmm can you post a quick screenshot of the material?

geerky42 (geerky42):

so "brutal force" is ok to you? lol

OpenStudy (anonymous):

so, just trials method ?

geerky42 (geerky42):

to my limited knowledge, yeah, though I'm certain there is better method.

OpenStudy (anonymous):

btw, 7 is works too :)

geerky42 (geerky42):

@amistre64 Do you know any methods?

geerky42 (geerky42):

Do you just need one value of p that works or all possible values of p?

OpenStudy (anonymous):

all values

geerky42 (geerky42):

then yeah brutal force wouldn't work here, since obviously there are infinity values of p we need to check lol...

geerky42 (geerky42):

@dan815 @Michele_Laino do you know any methods?

OpenStudy (michele_laino):

I try!

OpenStudy (michele_laino):

please wait...

OpenStudy (michele_laino):

I have found one solution.

OpenStudy (michele_laino):

I rewrite your equation as below: \[\large {x^3} - 4x + 9 = 0\]

geerky42 (geerky42):

Why set it equal to 0?

OpenStudy (michele_laino):

sorry I have misunderstood!

geerky42 (geerky42):

@texaschic101 @tkhunny @jim_thompson5910

OpenStudy (amistre64):

\[p^3-4p+9=x^2\] \[p(p^2-4)=x^2-9\] \[p(p-2)(p+2)=(x-3)(x+3)\] \[p=\frac{(x-3)(x+3)}{(p-2)(p+2)}\] \[\frac ab\in Z~\implies~\forall (a,b)\in Z,~and~a\ge b~:~gcd(a,b)=b\] can we work with this any?

OpenStudy (amistre64):

\[x^2-9~\ge p^2-4\] \[x^2-5~\ge p^2\] \[x^2-5~\ge p^2\] ------------------ \[x^2-9= k(p^2-4)\] \[\frac{x^2-9-4k}{k}= p^2\] ------------------------ \[kx^2-5k\ge x^2-4k-9\] \[(k-1)x^2-k+9\ge 0\] can we do anything with this nonesnse?